(a) A solid sphere, made of an insulating material, has a volume charge density of p , where r is the radius from the center of the sphere, a is constant, and a >0. What is the electric field within the sphere as a function of the radius r? Note: The volume element dv for a spherical shell of radius r and thickness dr is equal to 4tr2dr. (Use the following as necessary: a, r, and co.) magnitude E direction radially outward (b) what If? what if the charge density as a function of r within the charged solid sphere is given by ρ direction of the electric field within the sphere at radius r. (Use the following as necessary: a, r, and ) magnitudeE direction Find the new magnitude and radially outward # ) .
Answer #1

a) From Gauss's law,

fo f Ε.ds-foE *4m근-q έρ where q' is the charge enclosed by the sphere of radius r.

Now -IpW- 1a /r.tradr ,. 2nd . Therefore E=2pi ar^{2}/4pi epsilon _{0}r^{2}=a/2epsilon _{0} .

This is the electric field within the sphere as a function of radius r.

b)In this case ho =a/r^{2}.

Thus q'=int ho dV=int a/r^{2}*4pi r^{2}dr=4pi ar .

Therefore E=4pi ar/4pi epsilon _{0}r^{2}=a/epsilon _{0}r .